OpenStudy (anonymous):

in order to use the elimination method for the x terms the x terms must add to equal 0 explain the steps you would take to solve the system by eliminating the x term

OpenStudy (cwrw238):

take an example of a system of equations in x and y which I have labelled ! and 2: x + y = 7.......................(1) 2x - 3y = -1 .....................(2) we need to eliminate x so the first step is to multiply (1) by -2:- -2x - 2y = -14 .....................(3) 2x - 3y = -1 .....................(2) now adding (3) and (2) will eliminate x because -2x + 2x = 0 adding we get 0 - 5y = -15 we can now find the value of y

OpenStudy (cwrw238):

y = -15/-5 = 3 to find x we plug in the value y = 3 in equation (1):- x + 3 = 7 x = 7-3 = 4 so the solution is x=4 , y = 3

OpenStudy (anonymous):

i dont get the second step

OpenStudy (cwrw238):

y = -15/-5 = 3?

OpenStudy (cwrw238):

or adding the 2 equations (3) and (2) ?